OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)? 1) y = 4x - 2 2) y = 2x - 4 3) y = -3(x - 4) 4) y = 3(x + 4)

OpenStudy (anonymous):

i think it's 2 but im not sure

OpenStudy (sidsiddhartha):

equation of a line passing through (x1,y1) and (x2,y2) is \[\frac{ x-x_1 }{ y-y_1 }=\frac{ x_1-x_2 }{ y_1-y_2}\]

OpenStudy (sidsiddhartha):

now put the values of x1,y1 and x2,y2 and u'll get a equation

OpenStudy (anonymous):

4,2 over 0,6?

OpenStudy (anonymous):

wait what? im so confused right now

OpenStudy (sidsiddhartha):

 so x1=4 y1=0 x2=2 y2=6 now put them in the above equation